if the diameter of a wheel is 2 feet.approximately how many times it rotates to cross 2 miles?
Answer by mananth(16946)

diameter = 2feet
cicumference = pi*d
=3.142*2
=6.284 feet.
when the wheel makes one rotation it travels 6.284 feet.
it has to travel 2 miles
1 mile = 5280 feet
2 miles = 10560 feet
number of rotations = 10560/6.284
= 1680 rotations
...
m.ananth@hotmail.ca